in an evacuated 2.0 liter container, 1.8 moles of N2 and 2.2 moles of H2 are placed. at equilibrium 0.6 M of NH3 are present. solve for Kc
1
Expert's answer
2020-11-09T14:06:18-0500

"\begin{aligned}
N_2 + 3H_2 \to 2NH_3
\end{aligned}"


Volume of cylinder = 2.0litre = 2.0"dm^3"

Concentration of "N_2" = 1.8moles/2 "dm^3" = 0.9M

Concentration of "H_2" = 2.2moles/2 "dm^3" = 1.10M


Before reaction;

0.9(N) + 1.10(H) = 0(NH3)

During reaction

-x(N) -3x(H) = 2x(NH3)

After reaction;

0.9-x(N) + 1.10-3x(H) = 0.6(NH3)


"\therefore" x = 0.3M

Therefore, final conc. of "N_2= 0.9-0.3= 0.6M"

Final conc. of "H_2= 1.10-3(0.3) = 0.2M"


"K_c= \dfrac{[NH_3]^2}{[N_2][H_2]^3}"


"K_c= \dfrac{[0.6]^2}{[0.6][0.2]^3}"


"K_c = 75"